# Why Machine Learning Models Systematically Underestimate Extreme Values

### v1

@@ 2.4 Theoretical Framework - Multivariate Linear Regression
where $\beta_{j}$ are the true coefficients and $\delta_{y}$ represents the combined measurement uncertainty and intrinsic scatter in $y$, with $\mathbb{E}[\delta_{y}]=0$ and $\text{Var}(\delta_{y})=\sigma^{2}_{y}$. As established earlier in the univariate case, we need not distinguish between these sources of uncertainty in $y$ as they do not affect the attenuation bias.
@@
Substituting $\mathbf{X}_{\text{obs}}=\mathbf{X}_{\text{true}}+\boldsymbol{\Delta}_{x}$ into Equation (26):
$\mathbb{E}[\hat{\boldsymbol{\beta}}]$ $=\mathbb{E}[(\mathbf{X}_{\text{obs}}^{\top}\mathbf{X}_{\text{obs}})^{-1}\mathbf{X}_{\text{obs}}^{\top}\mathbf{y}_{\text{o}bs}]$ (27)
@@
$\mathbb{E}[\mathbf{X}_{\text{obs}}^{\top}\mathbf{y}]=\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]\boldsymbol{\beta}$ (32)
@@ 2.4.1 Case 1: Independent Multivariate Features
Let’s first examine the case where the features $x_{\text{true},j}$ are mutually independent. This scenario commonly occurs when different factors contribute additively to an outcome without interacting with each other. Consider, for example, a student’s exam score, which might depend on separate factors like hours of study, quality of sleep, and completion of practice tests—each independently affecting the final result. However, we should note that this example, while illustrative, rarely applies directly to astronomical analysis, as astronomical features $\mathbf{x}$ typically exhibit strong interdependencies. We will explore these more general and realistic cases in the next subsection.
When features are independent, $\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]$ simplifies to a diagonal matrix where each diagonal element represents the variance of the corresponding feature multiplied by the sample size:
@@
This result reveals a crucial insight: for independent features of equal strength, the attenuation bias is independent of the number of features $p$. In other words, the bias induced by measurement uncertainties in the input features cannot be mitigated simply by including more features in the model. The magnitude of attenuation remains constant regardless of the dimensionality of the input feature space, maintaining the same form we derived in the univariate case.
@@ 2.4.2 Case 2: Correlated Multivariate Features
The derivations above paint a bleak prospect for mitigating attenuation bias, as they apply to the case where $\mathbf{x}$ causes $y$ and all dimensions of the input $\mathbf{x}$ contribute to $y$ independently, like the exam score example laid out earlier. However, in most astronomical applications, there is still a saving grace stemming from the fact that it is the label $y$, the physical property of the object, that generates all the features $\mathbf{x}$, i.e., $y$ causes $\mathbf{x}$. By definition, in many astronomical applications, the features $\mathbf{x}$ are not independent (their measurement errors can be independent, but not $\mathbf{x}_{\text{true}}$ themselves). As we will demonstrate below, such correlation in features helps mitigate the attenuation bias as the number of features increases.
@@
$\mathbb{E}[\hat{\boldsymbol{\beta}}]\approx\lambda_{p}\mathbf{\beta},$ (43)
where the attenuation factor $\lambda_{p}$ is:
$\lambda_{p}=\frac{\sum_{j=1}^{p}a_{j}^{2}}{\frac{n\sigma_{x}^{2}}{S_{x}}+\sum_{j=1}^{p}a_{j}^{2}}.$ (44)
@@
$\lambda_{p}=\frac{1}{1+\frac{1}{\sigma^{\prime 2}}}.$ (46)
Interestingly, unlike the independent feature case, for a large number of correlated and non-trivial features ($p\to\infty$), we have $\sigma^{\prime}\to\infty$, and the attenuation factor $\lambda_{p}$ approaches 1, indicating that the bias can be mitigated by increasing the dimensionality of the input space, provided the features are sufficiently correlated. The attenuation bias in the correlated case is less severe than in the independent case, as the correlation among features allows the model to leverage additional information to reduce the impact of measurement errors.
Figure 4: Effect of dimensionality on attenuation bias in coefficients and predictions for independently drawn input features in multivariate linear regression. Left panel shows the distribution of coefficient attenuation factors ($\lambda_{\beta}=\hat{\beta}/\beta_{\text{true}}$), pooling results across all coefficients and simulations for each $p$, with boxcars showing the distribution over all coefficients and 100 simulations. Right panel shows prediction attenuation factors ($\lambda_{y}$), obtained by regressing predicted values against true values across all samples in each simulation, with boxcars showing the distribution over all simulations. For both panels, we examined three different signal-variance-to-noise ratios ($\sigma_{\text{range}}/\sigma_{x}=2,3,10$). The boxes indicate the interquartile range (IQR) with the median shown as a horizontal line, and whiskers extend to points within 1.5×IQR. The dashed lines show the theoretical predictions of $\lambda=1/(1+(\sigma_{x}/\sigma_{\text{range}})^{2})$ for each signal-variance-to-noise ratio. The inset plots show the variance of the coefficient estimates (left) and predicted values (right) as a function of dimensionality, with points showing empirical results and dashed lines showing theoretical predictions. The consistency between the left and right panels empirically confirms the theoretical derivation that the attenuation in coefficients directly translates to the same attenuation in predictions when measurement errors are homogeneous across dimensions, while the stability of attenuation factors across different $p$ values demonstrates that increasing the number of input dimensions does not mitigate the bias.
Figure 5: The lack of impact of training sample size on prediction attenuations for different ratios of signal range to measurement error with fixed input dimension $p=5$ for the multivariate linear regression with independent input features. Solid lines show mean attenuation factors ($\lambda_{y}$) with shaded regions indicating ±$\pm 1$ standard deviation calculated from 100 independent simulations. Results are shown for three ratios ($\sigma_{\text{range}}/\sigma_{x}=2,3,5$). For each ratio, the horizontal dashed line indicates its theoretical attenuation limit $\lambda=1/(1+(\sigma_{x}/\sigma_{\text{range}})^{2})$, while the curved dashed lines show the $1/\sqrt{n}$ convergence scaling arising from the variance of coefficient estimates. The results demonstrate how estimates become more precise with increasing sample size but remain fundamentally limited by the ratio of signal range to measurement error in the input features. The solid black line at $\lambda_{y}=1$ indicates the case of no attenuation.
2.5 Numerical Validations – Multivariate Linear Regression
@@ 2.5.1 Case 1: Independent Multivariate Features
To validate the theoretical framework of multivariate attenuation bias, we conducted numerical simulations examining both coefficient and prediction attenuation across varying numbers of input dimensions. We will first focus on the case where the features independently contribute to the output label.
@@
Figure 6: Effect of dimensionality on attenuation bias for perfectly correlated features, demonstrating mitigation of bias through feature correlation. Left panel shows coefficient attenuation factors ($\lambda_{\beta}$) and right panel shows prediction attenuation factors ($\lambda_{y}$) for different numbers of perfectly correlated input dimensions. Results are shown for three ratios of signal range to measurement error ($\sigma_{\text{range}}/\sigma_{x}=2,3,10$) based on 100 independent simulations. Features are generated as linear transformations of a uniform random variable on [0,1] ($\sigma_{\text{range}}=1/\sqrt{12}$) with scaling factors $a_{j}$ varying linearly from 0.2 to 0.4 across dimensions, chosen to ensure labels remain of order unity. In the left panel, box plots show the distribution of coefficient attenuation across all coefficients and simulations, with boxes offset horizontally for visibility across different $\sigma_{\text{range}}/\sigma_{x}$ ratios. The right panel shows the mean prediction attenuation with ±$\pm 1$ standard deviation across simulations. Dashed lines in both panels show theoretical predictions, demonstrating excellent agreement with the numerical results. Unlike the case with independent features (Figure 4), increasing dimensionality helps mitigate the attenuation bias. The solid black line at $\lambda=1$ indicates no attenuation.
As discussed in Section 2.4.1, in this simplifying limit where features independently contribute to the label, the attenuation factor for the coefficients $\lambda_{\beta}$ is analytically predictable. Furthermore, when both the coefficients $\beta_{j}$ and the features $\mathbf{x}$ are homogeneous across all pixels, we derived that the attenuation factor for predictions $\lambda_{y}$ should exactly follow $\lambda_{\beta}$. A key prediction from the theoretical framework is that such attenuation cannot be mitigated simply by increasing the number of pixels.
@@
In the case of independent features, not only does increasing the number of pixels fail to mitigate the attenuation bias, it actually comes with a cost. As we derived in Equation 37, the variance of the coefficients increases with $p$ due to the $\sum_{k=1}^{p}\beta_{k}^{2}$ term in the numerator. This behavior is perhaps not surprising from a geometric perspective: as dimensionality increases while the number of data points remains constant, we encounter the curse of dimensionality, where estimation of the hyperplane becomes increasingly stochastic. The variance of these linear coefficients follows a predictable pattern that we can derive analytically. Already seen in the boxcar plot, the spread increases with $p$. We further calculate the variance of these spreads. The inset plots in both panels reveal the variance behavior of our estimates as a function of dimensionality. As shown in the inset plot of the left panel, the empirical variance of coefficient estimates (points) closely follows the theoretical prediction (dashed lines) from Equation 37, demonstrating how the uncertainty scales systematically with the number of dimensions.
In the case where the label is simply the linear weighted sum of these variables, the increase in coefficient variance naturally leads to larger variance in the predictions. In the limit where the variance of $y=\boldsymbol{\beta}\cdot\mathbf{x}$ is dominated by the variance of $\boldsymbol{\beta}$, the sum of these random variables is still just a convolution of the various Gaussians, and thus larger variance in $\boldsymbol{\beta}$ translates into larger variance in the predicted $y$. However, as shown in the boxcar plot on the right panel, the prediction of $\lambda_{y}$ itself remains stable over $p$, since here we are taking the average over all pixels instead of the sum. In the inset plot of the right panel, the variance of $y_{\text{pred}}$ increases with $p$. We note that, to be conservative, we only calculate the variance of $y_{\text{pred}}$, since we know that $y_{\text{pred}}$ also suffers from an attenuation bias, and hence the variance of $y_{\text{pred}}-y_{\text{true}}$ would have a dependency on $\sigma_{\text{range}}/\sigma_{x}$, but the variance of $y_{\text{pred}}$ itself does not.
@@
The results reveal that as sample size increases, the attenuation factor for each $\sigma_{\text{range}}/\sigma_{x}$ ratio converges to its respective theoretical value at a rate that follows a $1/\sqrt{n}$ scaling law (shown as dashed lines). This scaling behavior directly emerges from the variance of $\hat{\boldsymbol{\beta}}$ in our theoretical derivation and propagates to the predictions through the linear relationship. However, the key finding is that at all training sizes, we observe persistent attenuation in predictions that depends on the ratio $\sigma_{\text{range}}/\sigma_{x}$. Even in cases with larger sample sizes, predictions asymptotically approach their theoretical attenuation limits.
In summary, these results do highlight a fundamental limitation in multivariate regression with independent features: the systematic bias introduced by measurement uncertainties cannot be overcome simply by increasing the number of input dimensions or size of the training sample.
@@ 2.5.2 Case 2: Correlated Multivariate Features
We have established that deriving the analytic attenuation factor is challenging, if not impossible, for general cases for correlated features. To make the theoretical analysis tractable, as assumed in the theoretical framework, we set the regression coefficients $\boldsymbol{\beta}$ equal to the scaling factors $\mathbf{a}$, allowing us to derive analytical predictions for the attenuation factor of coefficients as shown in Equation 45 and 46. If we further assume homogeneous and identical (but not independent) $x_{\text{true}}$, the attenuation factor $\lambda_{\beta}$ then translates into the attenuation factor of the label as we take the expectation through the linear transformation. With this framework, we can check with simulations to build insight but to ensure also that the theoretical framework is accurate.
@@ 3 Attenuation Bias in Spectral Analysis
Such bias has been broadly observed in many machine learning tasks applied to APOGEE spectra (see for example fig 7-8 in Fabbro et al. 2018 or fig 9-11 in Leung & Bovy 2019). While much of the consensus often attributes these effects to the precision of training labels or sample size limitations, the theoretical framework and numerical validation show that these factors have no impact on such attenuation bias.
@@
Figure 9: Impact of correlated features on attenuation bias in [Fe/H] estimation from APOGEE spectra. We select pixels in order of decreasing sensitivity to [Fe/H]. The solid black line shows the theoretical prediction for the most sensitive pixel ($\lambda=1/(1+(\sigma_{x}/\sigma_{\text{range,max}})^{2})$), with yellow points showing empirical results using only this pixel. Points show attenuation factors ($\lambda_{y}$) for different numbers of input features, shown in yellow (1 pixel), green (3 pixels), red (10 pixels), and blue (100 pixels), with error bars indicating $\pm$1 standard deviation across 100 independent trials. Bottom $x$-axis shows measurement uncertainty ($\sigma_{x}$) in normalized flux units, while top $x$-axis shows the corresponding signal-to-noise ratio. The panels on top show predicted versus true [Fe/H] for both 10-pixel (top row, red) and 100-pixel (bottom row, blue) cases at selected $\sigma_{x}$ values. Even at SNR = 100, the 10-pixel case shows visible attenuation, which becomes even more obvious at SNR = 10. The improvement in attenuation factors with increasing number of pixels demonstrates the advantage of correlated features over the independent case, though the effect remains non-negligible in most spectral analysis scenarios, particularly for elemental abundances that contribute to only a finite number of spectral features and/or at low SNR.
@@ 3.1 Numerical Investigation of Attenuation Bias Using Synthetic APOGEE Spectra
Figure 10: Impact of stellar parameter variations on pixel correlations and attenuation bias. Left panel shows the distribution of pairwise correlation coefficients among the 100 most [Fe/H]-sensitive pixels for different ranges of temperature variation, calculated using kernel density estimation (bandwidth = 0.03). Vertical lines indicate mean correlations for each temperature range. With fixed temperature (orange), all pixels are perfectly correlated by construction. As temperature variations increase from $T_{\text{eff}}\pm 200$K (blue) to $T_{\text{eff}}\pm 500$K (green) and $T_{\text{eff}}\pm 2000$K (purple), correlations weaken and their distributions broaden. Right panel shows how these weakened correlations affect attenuation bias in the 100-pixel case, with points showing mean attenuation factors and error bars indicating $\pm$1 standard deviation across 100 independent trials. Bottom $x$-axis shows measurement uncertainty ($\sigma_{x}$) in normalized flux units, while top $x$-axis shows the corresponding signal-to-noise ratio assuming $\sigma_{\text{range}}=0.1$. The increasing temperature variations progressively diminish the protective effect of correlated features.
@@
For each temperature range, we calculate the correlation matrix for our 100 most sensitive pixels. Taking the upper triangular portion of this matrix to avoid redundancy, we compute the distribution of unique pairwise correlation coefficients. The left panel shows these distributions using kernel density estimation (KDE with bandwidth 0.03), with vertical lines indicating the mean correlation for each temperature range. By definition with zero temperature spread, we only have correlation of one which is shown as the solid vertical line on right panel. As expected, increasing temperature variation progressively weakens pixel correlations. With no temperature variation (orange), the pixels show nearly perfect correlation as expected from our previous analysis. However, as we increase the temperature range to $T_{\text{eff}}\pm 200$K (blue), $T_{\text{eff}}\pm 500$K (green), and $T_{\text{eff}}\pm 2000$K (purple), the correlation distributions broaden and shift toward lower values even for these most sensitive pixels responding to [Fe/H]. This decorrelation occurs because different spectral features respond differently to temperature changes—some features become stronger while others weaken, even at the same [Fe/H].

### v2

@@ 2.4 Theoretical Framework - Multivariate Linear Regression
where $\beta_{j}$ are the true coefficients and $\delta_{y}$ represents the combined measurement uncertainty and intrinsic scatter in $y$, with $\mathbb{E}[\delta_{y}]=0$ and $\text{Var}(\delta_{y})=\sigma^{2}_{y}$. Since uncertainties in $y$ do not interact with the independent variables, they have identical effects on the regression analysis and do not influence the attenuation bias, as demonstrated in our univariate case.
@@
$\mathbb{E}[\mathbf{X}_{\text{obs}}^{\top}\mathbf{y}_{\text{obs}}]=\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]\boldsymbol{\beta}$ (31)
@@ 2.4.1 Case 1: Independent Multivariate Features
Let’s first examine the case where the features $x_{\text{true},j}$ are mutually independent. This scenario commonly occurs when different factors contribute additively to an outcome without interacting with each other. When features are independent, $\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]$ simplifies to a diagonal matrix where each diagonal element represents the variance of the corresponding feature multiplied by the sample size:
@@
This result reveals that, for independent features of equal strength, the attenuation bias is independent of the number of features $p$. The bias induced by measurement uncertainties in the input features cannot be mitigated simply by including more features in the model.
@@
Figure 3: Effect of dimensionality on attenuation bias in coefficients and predictions for independently drawn input features in multivariate linear regression. Left panel shows the distribution of coefficient attenuation factors ($\lambda_{\beta}=\hat{\beta}/\beta_{\text{true}}$), pooling results across all coefficients and simulations for each $p$, with boxcars showing the distribution over all coefficients and 100 simulations. Right panel shows prediction attenuation factors ($\lambda_{y}$), obtained by regressing predicted values against true values across all samples in each simulation, with boxcars showing the distribution over all simulations. For both panels, we examined three different signal-variance-to-noise ratios ($\sigma_{\text{range}}/\sigma_{x}=2,3,10$). The boxes indicate the interquartile range (IQR) with the median shown as a horizontal line, and whiskers extend to points within 1.5×IQR. The open circles show individual realizations that fall beyond the whiskers. The dashed lines show the theoretical predictions of $\lambda=1/(1+(\sigma_{x}/\sigma_{\text{range}})^{2})$ for each signal-variance-to-noise ratio. The inset plots show the variance of the coefficient estimates (left) and predicted values (right) as a function of dimensionality, with points showing empirical results and dashed lines showing theoretical predictions.
@@ 2.4.2 Case 2: Correlated Multivariate Features
In most astronomical applications, there is still a saving grace stemming from the fact that it is the label $y$, the physical property of the object, that generates all the features $\mathbf{x}$, i.e., $y$ causes $\mathbf{x}$. By definition, in many astronomical applications, the features $\mathbf{x}$ are not independent (their measurement errors can be independent, but not $\mathbf{x}_{\text{true}}$ themselves). As we will demonstrate below, such correlation in features helps mitigate the attenuation bias as the number of features increases.
@@
$\mathbb{E}[\hat{\boldsymbol{\beta}}]\approx\lambda_{\beta}\mathbf{\beta},$ (42)
where the attenuation factor $\lambda_{\beta}$ is:
$\lambda_{\beta}=\frac{\sum_{j=1}^{p}a_{j}^{2}}{\frac{n\sigma_{x}^{2}}{S_{x}}+\sum_{j=1}^{p}a_{j}^{2}}.$ (43)
@@
$\lambda_{\beta}=\frac{1}{1+\frac{1}{\sigma^{\prime 2}}}.$ (45)
Interestingly, unlike the independent feature case, for a large number of correlated and non-trivial features ($p\to\infty$), we have $\sigma^{\prime}\to\infty$, and the attenuation factor $\lambda_{\beta}$ approaches 1, indicating that the bias can be mitigated by increasing the dimensionality of the input space, provided the features are sufficiently correlated. The attenuation bias in the correlated case is less severe than in the independent case, as the correlation among features allows the model to leverage additional information to reduce the impact of measurement errors.
2.5 Numerical Demonstrations – Multivariate Linear Regression
@@ 2.5.1 Case 1: Independent Multivariate Features
To illustrate the theoretical framework of multivariate attenuation bias, we conducted numerical simulations examining both coefficient and prediction attenuation across varying numbers of input dimensions. We will first focus on the case where the features independently contribute to the output label.
@@
Figure 4: Prediction attenuations versus training sample size for different ratios of signal range to measurement error with fixed input dimension $p=5$ for multivariate linear regression with independent input features. Solid lines show mean attenuation factors ($\lambda_{y}$) with shaded regions indicating ±$\pm 1$ standard deviation calculated from 100 independent simulations. Results shown for three ratios ($\sigma_{\text{range}}/\sigma_{x}=2,3,5$). For each ratio, horizontal dashed line indicates theoretical attenuation limit $\lambda=1/(1+(\sigma_{x}/\sigma_{\text{range}})^{2})$, curved dashed lines show $1/\sqrt{n}$ convergence scaling from coefficient estimate variance. Solid black line at $\lambda_{y}=1$ indicates case of no attenuation.
Figure 5: Effect of dimensionality on attenuation bias for perfectly correlated features. Left panel shows coefficient attenuation factors ($\lambda_{\beta}$) and right panel shows prediction attenuation factors ($\lambda_{y}$) for different numbers of perfectly correlated input dimensions. Results are shown for three ratios of signal range to measurement error ($\sigma_{\text{range}}/\sigma_{x}=2,3,10$) based on 100 independent simulations. Features are generated as linear transformations of a uniform random variable on [0,1] ($\sigma_{\text{range}}=1/\sqrt{12}$) with scaling factors $a_{j}$ varying linearly from 0.2 to 0.4 across dimensions, chosen to ensure labels remain of order unity. In the left panel, box plots show the distribution of coefficient attenuation across all coefficients and simulations, with boxes offset horizontally for visibility across different $\sigma_{\text{range}}/\sigma_{x}$ ratios. The right panel shows the mean prediction attenuation with ±$\pm 1$ standard deviation across simulations. Dashed lines in both panels show theoretical predictions. The solid black line at $\lambda=1$ indicates no attenuation.
@@
In the case of independent features, not only does increasing the number of pixels fail to mitigate the attenuation bias, it actually comes with a cost. As we derived in Equation 36, the variance of the coefficients increases with $p$ due to the $\sum_{k=1}^{p}\beta_{k}^{2}$ term in the numerator. This behavior is perhaps not surprising from a geometric perspective: as dimensionality increases while the number of data points remains constant, we encounter the curse of dimensionality, where estimation of the hyperplane becomes increasingly stochastic. The inset plots in both panels reveal the variance behavior of our estimates as a function of dimensionality. As shown in the inset plot of the left panel, the empirical variance of coefficient estimates (points) closely follows the theoretical prediction (dashed lines) from Equation 36, demonstrating how the uncertainty scales systematically with the number of dimensions.
@@
The results reveal that as sample size increases, the attenuation factor for each $\sigma_{\text{range}}/\sigma_{x}$ ratio converges to its respective theoretical value at a rate that follows a $1/\sqrt{n}$ scaling law (shown as dashed lines). This scaling behavior directly emerges from the variance of $\hat{\boldsymbol{\beta}}$ in our theoretical derivation and propagates to the predictions through the linear relationship. However, the key finding is that at all training sizes, we observe persistent attenuation in predictions that depends on the ratio $\sigma_{\text{range}}/\sigma_{x}$. The systematic bias introduced by measurement uncertainties cannot be overcome simply by increasing the number of input dimensions or size of the training sample.
@@ 2.5.2 Case 2: Correlated Multivariate Features
We have established that deriving the analytic attenuation factor is challenging, if not impossible, for general cases for correlated features. To make the theoretical analysis tractable, as assumed in the theoretical framework, we set the regression coefficients $\boldsymbol{\beta}$ equal to the scaling factors $\mathbf{a}$, allowing us to derive analytical predictions for the attenuation factor of coefficients as shown in Equation 44 and 45. If we further assume homogeneous and identical (but not independent) $x_{\text{true}}$, the attenuation factor $\lambda_{\beta}$ then translates into the attenuation factor of the label as we take the expectation through the linear transformation. With this framework, we can check with simulations to build insight but to ensure also that the theoretical framework is accurate.
@@ 3 Attenuation Bias in Spectral Analysis
Such bias has been broadly observed in many machine learning tasks applied to APOGEE spectra (see for example fig 7-8 in Fabbro et al. 2018 or fig 9-11 in Leung and Bovy 2019). While much of the consensus often attributes these effects to the precision of training labels or sample size limitations, the theoretical framework and numerical validation show that these factors have no impact on such attenuation bias.
@@ 3.1 Numerical Investigation of Attenuation Bias Using Synthetic APOGEE Spectra
[…]
Figure 8: Impact of correlated features on attenuation bias in [Fe/H] estimation from APOGEE spectra. We select pixels in order of decreasing sensitivity to [Fe/H]. The solid black line shows the theoretical prediction for the most sensitive pixel ($\lambda=1/(1+(\sigma_{x}/\sigma_{\text{range,max}})^{2})$), with yellow points showing empirical results using only this pixel. Points show attenuation factors ($\lambda_{y}$) for different numbers of input features, shown in yellow (1 pixel), green (3 pixels), red (10 pixels), and blue (100 pixels), with error bars indicating $\pm$1 standard deviation across 100 independent trials. Bottom $x$-axis shows measurement uncertainty ($\sigma_{x}$) in normalized flux units, while top $x$-axis shows the corresponding signal-to-noise ratio. The panels on top show predicted versus true [Fe/H] for both 10-pixel (top row, red) and 100-pixel (bottom row, blue) cases at selected $\sigma_{x}$ values. Even at SNR = 100, the 10-pixel case shows visible attenuation, which becomes even more obvious at SNR = 10.
@@
Figure 9: Distribution of pairwise correlation coefficients among the 100 most [Fe/H]-sensitive pixels for different ranges of temperature variation, calculated using kernel density estimation (bandwidth = 0.03). Left panel: Vertical lines indicate mean correlations for each temperature range. Fixed temperature case shown in orange, with temperature variations of $T_{\text{eff}}\pm 200$K (blue), $T_{\text{eff}}\pm 500$K (green) and $T_{\text{eff}}\pm 2000$K (purple). Right panel: Mean attenuation factors and $\pm$1 standard deviation error bars across 100 independent trials for the 100-pixel case. Bottom $x$-axis shows measurement uncertainty ($\sigma_{x}$) in normalized flux units, top $x$-axis shows corresponding signal-to-noise ratio assuming $\sigma_{\text{range}}=0.1$.
@@
For each temperature range, we calculate the correlation matrix for our 100 most sensitive pixels. Taking the upper triangular portion of this matrix to avoid redundancy, we compute the distribution of unique pairwise correlation coefficients. The left panel shows these distributions using kernel density estimation (KDE with bandwidth 0.03), with vertical lines indicating the mean correlation for each temperature range. By definition with zero temperature spread, we only have correlation of one which is shown as the solid vertical line on right panel. As expected, increasing temperature variation progressively weakens pixel correlations. With no temperature variation (orange), the pixels show nearly perfect correlation as expected from our previous analysis.
However, as we increase the temperature range to $T_{\text{eff}}\pm 200$K (blue), $T_{\text{eff}}\pm 500$K (green), and $T_{\text{eff}}\pm 2000$K (purple), the correlation distributions broaden and shift toward lower values even for these most sensitive pixels responding to [Fe/H]. The ±$\pm 2000$K range represents an extreme case where temperature information is essentially not measured, allowing us to examine the limiting behavior of decorrelation. This decorrelation occurs because different spectral features respond differently to temperature changes—some features become stronger while others weaken, even at the same [Fe/H].